OpenStudy (anonymous):

The number of computers sold by BCC depends on the dollar amount, x, that they spend on advertising. How much must they spend to sell 200 computers? Round to the nearest dollar and do not include units in your answer. N(x) = 100 + 20 * ln(0.25x)

OpenStudy (anonymous):

set \[N(x)=200=100+20\times \ln(.25x)\] and solve for x

OpenStudy (anonymous):

\[200=100+20\times\ln(.25x)\] \[100=20\times \ln(.25x)\] \[5=\ln(.25x)\] \[.25x=e^5\] \[x=4e^5\]

OpenStudy (anonymous):

that looks better